How to find maximum value cosAcosBcosC in the plane triangle
 

Grade:12th pass

1 Answers

Arun
5 years ago
If any of A, B and C are not acute, then clearly the product is non-positive (since cosx ≤ 0 for π/2 ≤ x ≤ π, and at most one angle can be non-acute. Thus we have the product of 2 positive numbers and a non-positive number, which is thus non-positive) 

Thus if any of A, B, and C are not acute then 
cosA cosB cosC ≤ 0 

Now if they are all acute then we have that cosA, cosB and cosC are all positive. 
Thus, by AM-GM, we have that (cosA + cosB + cosC)/2 ≥ (cosA cosB cosC)^(1/3) 
But we know that cosA + cosB + cosC ≤ 3/2 by problem 1. 
Thus (cosA cosB cosC)^1/3 ≤ (3/2)/3 = 1/2 
Cubing each side we get 
cosA cosB cosC ≤ 1/8 
(We can do this since f(x) = x^3 is increasing) 

Again, we have equality if the triangle is equilateral. 
Thus the maximum value of cosA cosB cosC is 1/8